### v1

# Digital Quantum Simulation of the Nonlinear Lindblad Master Equation Based on Quantum Trajectory Averaging

Abstract
Since precisely controlling dissipation in realistic environments is challenging, digital simulation of the Lindblad master equation (LME) is of great significance for understanding nonequilibrium dynamics in open quantum systems. However, achieving long-time simulations for complex systems with multiple dissipation channels remains a major challenge, both theoretically and experimentally. Here, we propose a 2-dilation digital simulation scheme for the non-linear Lindblad master equation (NLME) based on quantum trajectory averaging. The NLME continuously interpolates between full LME and the dynamical equation governed by the effective non-Hermitian Hamiltonian. Remarkably, for standard LMEs, our scheme reduces to a 1-dilation method that enables deterministic realizations without postselection. This deterministic nature overcomes a key limitation in some existing simulation methods, where repeated postselections lead to exponentially vanishing implementation probabilities. Consequently, our scheme allows efficient long-time simulations of LMEs with multiple jump operators. As a demonstration, we present numerical experiments simulating novel theoretical predictions in open quantum systems, including localization in open quantum systems and the postselected skin effect.
Introduction.— The Lindblad master equation (LME) plays a crucial role in the study of open quantum systems, providing a rigorous framework to describe the quantum system coupling with the environment under the Born-Markov approximation Lindblad1976 ; Gardiner1985 ; Daniel2020 . In recent years, the demand for precise control over dissipative processes has driven extensive theoretical efforts to explore novel nonequilibrium dynamical phenomena within the Lindblad framework, such as Liouvillian flat bands Hugo2023 , postselected skin effects Hugo2025 , and localization in open quantum systems Yusipov2017 ; YLiu2024 . These studies often require carefully engineered dissipation operators, which are difficult to implement on analog simulation platforms. In contrast, digital quantum platforms provide a standardized and promising approach to simulation.
The primary challenge of digitally simulating the LME lies in implementing its intrinsically non-unitary nature using unitary logic gates. This difficulty can be addressed through approaches such as dilation methods RCleve2017 ; ZDing2024 ; ZHu2020 ; KMarsden2021 ; ZHu2022 ; ASchlimgen2021 ; ASchlimgen2022 ; NSuri2023 ; DMazziotti2022 ; EOh2024 ; SPeng2024 ; LHu2019 ; JHan2021 , which effectively embed non-unitarity into enlarged unitary gates, and hybrid quantum-classical algorithms, including variational quantum simulation XYuan2019 ; SEndo2020 ; Mahdian2020 ; JLuo2024 and the quantum imaginary time evolution method MMotta2020 ; HKamakari2022 . Currently, effective experimental schemes have been demonstrated for single-qubit and few-qubit systems LHu2019 ; JHan2021 ; ASchlimgen2021 ; ZHu2020 ; ZHu2022 ; EOh2024 ; HKamakari2022 ; JWPan2017 . However, for large systems, numerous dissipation sources, and long-time evolution, significant theoretical and experimental challenges remain. For instance, the Stinespring dilation approach RCleve2017 ; ZDing2024 requires a large number of auxiliary qubits, potentially affecting its scalability. Moreover, if probabilistic approaches, such as Sz.-Nagy dilation (SND) ZHu2020 ; KMarsden2021 ; ZHu2022 , the linear combination of unitaries (LCU) method ASchlimgen2021 ; ASchlimgen2022 ; NSuri2023 , and singular-value decomposition (SVD) method DMazziotti2022 ; EOh2024 are used to iteratively simulate system evolution within each time step $\delta t$, the overall success probability rapidly approaches zero as the number of time steps and dissipation sources increases.
An alternative simulation scheme for the LME on a classical computer is the quantum trajectory method Dalibard1992 ; Dum1992 ; Gardiner1992 ; Klaus1993 ; Plenio1998 ; Daley2014 ; HWiseman2016 , which interprets the LME as a stochastic average over individual trajectories of pure states. This approach significantly reduces memory usage, as it focuses on the state rather than the density matrix. Inspired by this method, Luo et al. JLuo2024 and Peng et al. SPeng2024 recently proposed new variational and dilation-based simulation schemes, which substantially reduce logical gate complexity. However, the former requires additional ansatz assumptions, while the latter remains a probabilistic implementation and still faces the challenge of an almost vanishing success probability at long times.
In our trajectory simulation scheme, we overcome the challenge of the overall success probability approaching zero. The key insight is to rigorously match the stochasticity in quantum trajectories with the probabilistic outcomes of measurements in quantum circuits. Each measurement outcome directly corresponds to a specific mathematical process in the quantum trajectory method: either a quantum jump or non-unitary evolution. Unlike methods based on SND, LCU, and SVD, where certain states are discarded after measurement, our approach ensures that every measurement outcome is meaningful.
Based on the principle of efficiently utilizing measurement randomness, our scheme can extend from deterministic trajectory simulations of the LME to probabilistic trajectory simulations of the non-linear Lindblad master equation (NLME) Hugo2025 ; Lewalle2023 ; MCoppola2024 ; APaviglianiti2024 . The NLME describes the dynamics of an open system lacking an arbitrary proportion of quantum jumps, providing a continuous interpolation between the LME and the evolution of effective non-Hermitian Hamiltonian (ENHH). It gives a reasonable framework for studying the interplay and competition between the non-Hermitian Hamiltonians and pure dissipation or pure measurement processes.
In this paper, we first introduce the trajectory-averaged interpretation of the NLME. Based on this interpretation, we propose a probabilistic 2-dilation implementation. For two special cases of the NLME, namely the LME and the ENHH, the scheme reduces to a 1-dilation. In particular, for the LME, the scheme becomes fully deterministic due to the absence of post-selection effects. Furthermore, we simulate a specific class of jump operators that are theoretically interesting but difficult to realize experimentally, demonstrating a numerical approach for simulating open-system localization and the postselected skin effect.
Figure 1: Scheme of the quantum trajectory simulation: (a) the 2-dilation method for the NLME as Eq. (1), (b) the 1-dilation method for the LME as Eq. (2), and (c) the 1-dilation method for the evolution of ENHH as Eq. (3). The system is initialized in state $|\phi(0)\rangle$. The auxiliary qubits are initialized in state $|00\rangle_{a}$ in (a) and in state $|0\rangle_{a}$ in both (b) and (c). At each time step $\delta t$, the operations inside the brackets are executed. The gate $U_{0}=e^{-iH\delta t}$ represents the evolution from system Hamiltonian. The gate $U_{\mu\geq 1}$, corresponding to the $\mu$-th dissipation source, is given by Eq. (7) in (a), Eq. (11) in (b), and Eq. (12) in (c). After each measurement of auxiliary qubit, the postselection, denoted by ’P’, is required only in (a) and (c), making their realization probabilistic. In contrast, the method in (b) is deterministic as it does not require postselection. In all three cases, the auxiliary qubit must be repeatedly reset to its initial state, denoted by ’R’.
Trajectory-averaged interpretation of NLME.— The NLME is a generation of the LME Hugo2025 ; Lewalle2023 . It continuously interpolates between LME and ENHH by reducing quantum jumps, given by
$\begin{split}&\frac{d}{dt}\rho=\mathcal{L}_{\eta}(\rho):=-i[H,\rho]+\sum_{\mu=1}^{M}\mathcal{D}_{\mu}(\rho),\\ &\mathcal{D}_{\mu}(\rho)=\gamma_{\mu}\left(-\frac{1}{2}\{L_{\mu}^{{\dagger}}L_{\mu},\rho\}+(1-\eta_{\mu})L_{\mu}\rho L_{\mu}^{{\dagger}}+\eta_{\mu}\langle L_{\mu}^{{\dagger}}L_{\mu}\rangle\rho\right),\end{split}$ (1)
where the system with Hamiltonian $H$ couples to the environment by $M$ jump operators $L_{\mu}$ with strength $\gamma_{\mu}$. The $\eta_{\mu}$ ($0\leq\eta_{\mu}\leq 1$) is the postselection strength that calibrates the reduction ratio of quantum jumps. When $\eta_{\mu}=0$ for all $\mu$, the NLME reduces to the standard LME,
$\frac{d\rho}{dt}=\mathcal{L}(\rho):=-i[H,\rho]+\sum_{\mu=1}^{M}\gamma_{\mu}\left(L_{\mu}\rho L_{\mu}^{{\dagger}}-\frac{1}{2}\{L_{\mu}^{{\dagger}}L_{\mu},\rho\}\right).$ (2)
Conversely, when $\eta_{\mu}=1$ for all $\mu$, the NLME simplifies to the evolution of ENHH,
$\frac{d}{dt}\rho=-iH_{\rm eff}\rho+i\rho H^{\dagger}_{\rm eff}+\sum_{\mu=1}^{M}\gamma_{\mu}\langle L_{\mu}^{{\dagger}}L_{\mu}\rangle\rho,$ (3)
where the ENHH takes the form $H_{\rm eff}=H-\frac{i}{2}\sum_{\mu=1}^{M}\gamma_{\mu}L_{\mu}^{{\dagger}}L_{\mu}.$
The NLME can be interpreted by the quantum trajectory method Hugo2025 . The expectation value of an arbitrary observable $\hat{O}$ is given by the average over many trajectories: $\langle\hat{O}\rangle=\mathrm{Tr}(\hat{O}\rho)=\lim_{K\to\infty}\frac{1}{K}\sum_{i=1}^{K}\langle\phi_{i}(t)|\,\hat{O}\,|\phi_{i}(t)\rangle$. The quantum state at time $t+\delta t$ for the $i$-th trajectory is obtained from the state at time $t$ through a stochastic evolution:
$\begin{split}&|\phi_{i}(t+\delta t)\rangle=e^{-iH\delta t}\prod_{\mu=1}^{M}\mathcal{U}_{\mu}(\delta t)\ |\phi_{i}(t)\rangle,\\ &\mathcal{U}_{\mu}(\delta t)=\left\{\begin{matrix}(1-\eta_{\mu})\delta p_{\mu}:&\mathcal{N}L_{\mu}\\ 1-(1-\eta_{\mu})\delta p_{\mu}:&\mathcal{N}\exp(-\frac{1}{2}\gamma_{\mu}L^{\dagger}_{\mu}L_{\mu}\delta t)\end{matrix}\right.\ ,\end{split}$ (4)
where $\mathcal{N}$ denotes normalization process and $\delta p_{\mu}=\gamma_{\mu}\delta t\langle\phi_{i}(t)|L^{\dagger}_{\mu}L_{\mu}|\phi_{i}(t)\rangle$. The dissipation effect $\mathcal{U}_{\mu}(\delta t)$ represents a stochastic process where the system either jumps to the state $\mathcal{N}L_{\mu}|\phi(t)\rangle$ with the probability of $(1-\eta_{\mu})\delta p_{\mu}$ or evolves to the state $\mathcal{N}\exp(-\frac{1}{2}\gamma_{\mu}L^{\dagger}_{\mu}L_{\mu}\delta t)|\phi(t)\rangle$.
Digital quantum simulation.— The key challenge in implementing an iterative simulation for a single quantum trajectory of NLME from Eq. (4) lies in how to realize the stochastic process $\mathcal{U}_{\mu}(\delta t)$. To address this within a quantum circuit, we assume a unitary gate $U_{\mu}$ that acts on the Hilbert space of the system along with two auxiliary qubits as follows:
$U_{\mu}\,|00\rangle_{a}|\phi\rangle=|00\rangle_{a}\,C_{\mu}|\phi\rangle+|01\rangle_{a}\,B_{\mu}|\phi\rangle+|10\rangle_{a}\,A_{\mu}|\phi\rangle,$ (5)
where the subscript $a$ represents auxiliary qubits. After measuring the auxiliary qubits, the system probabilistically collapses into the state of $\mathcal{N}X_{\mu}|\phi\rangle$ ($X$ = $A$, $B$, or $C$) with a probability of $\langle X_{\mu}^{\dagger}X_{\mu}\rangle:=\langle\phi|X_{\mu}^{\dagger}X_{\mu}|\phi\rangle$.
To match the measurement outcomes of Eq. (5) with the stochastic results of $\mathcal{U}_{\mu}(\delta t)$, we assume that $A_{\mu}\propto\exp(-\frac{1}{2}\gamma_{\mu}L^{\dagger}_{\mu}L_{\mu}\delta t)$ and $B_{\mu}\propto L_{\mu}$. To precisely match their probabilities, we discard the $C_{\mu}|\phi\rangle$ state by postselecting the auxiliary qubits to not be in the $|00\rangle_{a}$ state, and require renormalized probability distribution to satisfy
$\frac{\langle B_{\mu}^{\dagger}B_{\mu}\rangle}{\langle A_{\mu}^{\dagger}A_{\mu}\rangle+\langle B_{\mu}^{\dagger}B_{\mu}\rangle}=(1-\eta_{\mu})\delta p_{\mu}.$ (6)
By enforcing both the probability constraint Eq. (6) and the unitarity condition $U_{\mu}^{\dagger}U_{\mu}=\mathbb{I}$, we obtain a solution for $U_{\mu}$ as
$U_{\mu}=\left(\begin{matrix}C_{\mu}&B_{\mu}&\tilde{A}_{\mu}&0\\ B_{\mu}&-C_{\mu}&0&\tilde{A}_{\mu}\\ A_{\mu}&0&-C_{\mu}^{\dagger}&-B_{\mu}^{\dagger}\\ 0&A_{\mu}&-B_{\mu}^{\dagger}&C_{\mu}^{\dagger}\end{matrix}\right),$ (7)
where
$\begin{split}&A_{\mu}=\sqrt{1-\gamma_{\mu}\delta tL_{\mu}^{\dagger}L_{\mu}}\,,\ \ \tilde{A}_{\mu}=\sqrt{1-\gamma_{\mu}\delta tL_{\mu}L_{\mu}^{\dagger}}\,,\\ &B_{\mu}=\sqrt{(1-\eta_{\mu})\gamma_{\mu}\delta t}\,L_{\mu}\,,\ \ C_{\mu}=\sqrt{\eta_{\mu}\gamma_{\mu}\delta t}\,L_{\mu}\,,\end{split}$ (8)
and note that $\exp(-\frac{1}{2}\gamma_{\mu}L^{\dagger}_{\mu}L_{\mu}\delta t)=A_{\mu}+\mathcal{O}(\delta t)$.
The 2-dilation method for realizing the full trajectory is presented in Fig. 1(a). The system and auxiliary qubits are initialized in the state $|\phi_{i}(0)\rangle$ and $|00\rangle_{a}$. Then, N iterations are carried out for a time evolution of duration $T=N\delta t$. During each time step $\delta t$, the unitary evolution governed by the system Hamiltonian is encoded in the logic gate $U_{0}=e^{-iH\delta t}$. Each dissipation effect $\mathcal{U}_{\mu}(\delta t)$ is implemented by measuring and post-selecting the auxiliary qubits after applying the 2-dilation logic gate $U_{\mu}$ in Eq. (7), where only the measurement outcomes of $|01\rangle_{a}$ and $|10\rangle_{a}$ are considered valid. If the outcome is $|00\rangle_{a}$ the iteration is ended, and this trajectory is discarded. Following postselection, the auxiliary qubits are reset to $|00\rangle_{a}$ state in preparation for the next gate $U_{\mu+1}$.
Since the postselection requires that the measurement outcome $|00\rangle_{a}$ cannot appear throughout the entire trajectory, the success probability for a single trajectory, with M dissipation sources and $N\delta t$ time duration, is
$\begin{split}P&=\prod_{t=\delta t}^{N\delta t}\prod_{\mu=1}^{M}\left(1-\langle\phi(t)|C_{\mu}^{\dagger}C_{\mu}\phi(t)|\rangle\right)\\ &=\prod_{t=\delta t}^{N\delta t}\prod_{\mu=1}^{M}\left(1-\eta_{\mu}\gamma_{\mu}\delta t\langle\phi(t)|L_{\mu}^{\dagger}L_{\mu}|\phi(t)\rangle\right).\end{split}$ (9)
$P$ is roughly estimated as $(1-\bar{\eta}\bar{\gamma}\langle L_{\mu}^{\dagger}L_{\mu}\rangle\delta t)^{NM}$, where $\bar{\eta}=\frac{1}{M}\sum_{\mu=1}^{M}\eta_{\mu}$ and $\bar{\gamma}=\frac{1}{M}\sum_{\mu=1}^{M}\gamma_{\mu}$.
When $\eta_{\mu}=0$ for all $\mu$, the NLME reduces to the standard LME in Eq. (2). At the same time, $C_{\mu}=0$ and postselection success probability reaches unity, i.e., $P=1$ in Eq. (9), which indicates that the 2-dilation probabilistic implementation of the NLME can reduce to the 1-dilation deterministic implementation of the LME, as shown in Fig. 1(b). Note that it does not require postselection because the stochasticity in $\mathcal{U}_{\mu}(t)$ is completely equivalent to the randomness of measurement outcomes. This means that after measuring the auxiliary qubit in the state
$U_{\mu}\,|0\rangle_{a}|\phi\rangle=|0\rangle_{a}\,B_{\mu}|\phi\rangle+|1\rangle_{a}\,A_{\mu}|\phi\rangle,$ (10)
where $U_{\mu}$ reduces to
$U_{\mu}=\left(\begin{matrix}B_{\mu}&\tilde{A}_{\mu}\\ A_{\mu}&-B_{\mu}^{\dagger}\end{matrix}\right),$ (11)
if the outcome is $|1\rangle_{a}$, the system has an evolvtion of $\mathcal{N}\exp(-\frac{1}{2}\gamma_{\mu}L^{\dagger}_{\mu}L_{\mu}\delta t)$ by the $A_{\mu}$, whereas the outcome of $|0\rangle_{a}$ corresponds to a quantum jump on the trajectory by the $B_{\mu}$. The consistency of the probabilities between $\mathcal{U}_{\mu}$ and measurement outcomes is guaranteed by the constraint equation Eq. (6).
When $\eta_{\mu}=1$ for all $\mu$, the 2-dilation method for the NLME can also reduce to the 1-dilation method for the ENHH in Eq. (3) due to $B_{\mu}=0$, as shown in Fig. 1(c). The gate $U_{\mu}$ and its action reduce to
$U_{\mu}\,|0\rangle_{a}|\phi\rangle=|0\rangle_{a}\,C_{\mu}|\phi\rangle+|1\rangle_{a}\,A_{\mu}|\phi\rangle;\ \ U_{\mu}=\left(\begin{matrix}C_{\mu}&\tilde{A}_{\mu}\\ A_{\mu}&-C_{\mu}^{\dagger}\end{matrix}\right).$ (12)
The postselection discards trajectories where the auxiliary qubit collapses to $|0\rangle_{a}$ after measurement.
Numerical experiments.— We first simulate a two-level atom with monitoring its spontaneous emission Hugo2025 ; SuMa to demonstrate the validity of our scheme. The dynamic equation is
${}\begin{split}\frac{d}{dt}\rho=&-iJ[\sigma^{x},\rho]\\ &+\gamma\left(-\frac{1}{2}\{\sigma^{+}\sigma^{-},\rho\}+(1-\eta)\sigma^{-}\rho\sigma^{+}+\eta\langle\sigma^{+}\sigma^{-}\rangle\rho\right),\end{split}$ (13)
where $\sigma^{\pm}$ is the spin raising and lowering operators and $\sigma^{x}=\sigma^{+}+\sigma^{-}$. This NLME can be exactly solved by the vectorization method Hugo2025 ; SuMa . We compare the evolution of excited-state probability from the exact solution with that from the digital simulation of our 2-dilation and 1-dilation methods in Fig. 1, where the curves exhibit good consistency.
[…]
The fundamental mission of quantum simulation is to explore complex systems. Now, we focus on a many-particle problem in an L-site open chain, with the evolution equation given by:
$\begin{split}~{}&\frac{d}{dt}\rho=-i[H,\rho]+\gamma\sum_{l=1}^{L-1}\mathcal{D}_{l}(\rho),\\ &H=J\sum_{l=1}^{L-1}\left(\sigma_{l}^{+}\sigma_{l+1}^{-}+{\rm H.c.}\right)+\sum_{l=1}^{L}V\cos(2\pi\omega l)\sigma_{l}^{z},\\ &\mathcal{D}_{l}(\rho)=-\frac{1}{2}\{L_{l}^{{\dagger}}L_{l},\rho\}+(1-\eta)L_{l}\rho L_{l}^{{\dagger}}+\eta\langle L_{l}^{{\dagger}}L_{l}\rangle\rho,\end{split}$ (14)
where $\sigma_{l}^{\mu}(\mu=z,+,-)$ is the spin-1/2 operator at the site $l$, and $\omega=(\sqrt{5}-1)/2$ indicates the quasicrystal potential. The local dissipation operator takes the following form:
${}L_{l}=\frac{1}{2}(\sigma_{l}^{+}+e^{i\alpha}\sigma_{l+1}^{+})(\sigma_{l}^{-}+e^{i\beta}\sigma_{l+1}^{-}),$ (15)
which represents the phase-changing effect between neighboring spins. This kind of dissipation operator exhibits novel phenomena in theoretical studies, such as entangled state preparation BKraus2008 ; SDiel2008 ; PSchindler2013 , localization Yusipov2017 ; YLiu2024 , and the skin effect in open quantum systems Hugo2025 ; Wang2022 ; Feng2023 ; ZCLiu2024 . Note that this model is an intrinsic many-body evolution due to the $\hat{n}_{l}\,\hat{n}_{l+1}$ interaction ($\hat{n}_{l}:=\sigma_{l}^{+}\sigma_{l}^{-}$) existing in the evolution factor of $-\frac{1}{2}\gamma L_{l}^{\dagger}L_{l}\delta t$.
[…]
Furthermore, when $\alpha=\pi/2$, $\beta=-\pi/2$, $V=0$, and $\eta\neq 0$, the steady state is expected to exhibit skin effect Hugo2025 . We simulate this postselected skin effect by our 2-dilation method in the supplementary materials SuMa .
Summary.— We tackle the scalability challenge in numerically simulating the LME over long-time evolutions with multiple dissipation sources. Based on quantum trajectory averaging, we propose a digital quantum simulation scheme for dynamics governed by the LME, ENHH, and NLME. The simulation of the NLME is achieved via a 2-dilation method, which simplifies to 1-dilation for the LME and ENHH. Notably, our approach enables deterministic simulation of the standard LME without postselection, distinguishing it from existing probabilistic schemes and allowing for long-time simulations of complex open systems.
Our work paves the way for efficient digital quantum simulations of the LME, enables the exploration of the interplay between non-Hermitian Hamiltonians and pure dissipation processes through NLME simulations, and offers a promising simulation scheme for various cutting-edge theoretical models in open quantum systems and non-Hermitian physics.
Acknowledgements.
Acknowledgments.— Y.-G. Liu thanks Zongsheng Zhou for the valuable discussions during the early stages of the work. The work is supported by National Key Research and Development Program of China (Grant Nos. 2021YFA1402104 and 2023YFA1406704), the National Natural Science Foundation of China (Grants Nos. 12474287, 12174436, T2121001, 92265207).